OpenStudy (anonymous):

Factor the trinomial: x2 - 14x + 24 (x - 3)(x - 8) (x - 2)(x - 12) (x - 4)(x + 12) prime

OpenStudy (anonymous):

(x-2)(x-12) is the answer

OpenStudy (amistre64):

test/ multiple choices are a different procedure than actually knowing what to do :) but same clues ... x2 - 14x (+) 24; that + tells us we add same signs x2 (-) 14x + 24; that - tells us that we need both to be - -8-3 = -11 -2-12 = -14 <- this one works